the ice skater rotates around its axis, arms outstretched, at an angular velocity "w1=4 rad\/s" . what will be the angular velocity if the slider gathers its hands and at the same time reduces the moment of inertia three times? the system is considered isolated.


1
Expert's answer
2021-03-04T17:24:34-0500

Let's assume the angular momentum of the system is conserved. Thus, we have:


"I_1\omega_1 = I_2\omega_2"

where "I_1 = 3I_2" are moments of inertia before and after arms gathering respectively, "\omega_1 = 4rad/s, \omega_2" are the angular velocities before and after arms gathering respectively. Thus, obtain:


"\omega_2 = \dfrac{I_1\omega_1}{I_2}=\dfrac{3I_2\omega_1}{I_2} = 3\omega_1\\
\omega_2 = 3\cdot 4rad/s = 12rad/s"

Answer. 12 rad/s.